OpenStudy (anonymous):

Find the surface area and volume of a sphere having a radius of 4. (Leave answer as an improper fraction.)

OpenStudy (anonymous):

surface area of a sphere=4*pi*r*r volume of the sphere=4/3*pi*r*r*r where r is the radius....